Use the definition of limit to prove that the sequence { 1/√n+1 } converges to 0
1
Expert's answer
2021-01-26T15:09:23-0500

The sequence "\left(\frac{1}{\sqrt{n}+1}\right)_n" converges to 0 if


"\forall \varepsilon>0,\;\exists N\in\mathbb{N},\; \left(n\geq N\implies \left|\tfrac{1}{\sqrt{n}+1}-0\right|=\tfrac{1}{\sqrt{n}+1}\leq\varepsilon\right)."


Let "\varepsilon>0". If we take "N=\left\lceil\dfrac{1}{\varepsilon}\right\rceil^2" we obtain "\sqrt{N}=\left\lceil\dfrac{1}{\varepsilon}\right\rceil\geq\dfrac{1}{\varepsilon}". Then "\dfrac{1}{\sqrt{N}}\leq\varepsilon". Therefore

"\frac{1}{\sqrt{N}+1}\leq\frac{1}{\sqrt{N}}\leq\varepsilon."

So, for all "n\geq N,"

"\frac{1}{\sqrt{n}+1}\leq\frac{1}{\sqrt{N}+1}\leq\epsilon"